# Understanding Knowledge Hijack Mechanism in In-context Learning through Associative Memory

### v1

@@ Abstract
In-context learning (ICL) enables large language models (LLMs) to adapt to new tasks without fine-tuning by leveraging contextual information provided within a prompt. However, ICL relies not only on contextual clues but also on the global knowledge acquired during pretraining for the next token prediction. Analyzing this process has been challenging due to the complex computational circuitry of LLMs. This paper investigates the balance between in-context information and pretrained bigram knowledge in token prediction, focusing on the induction head mechanism, a key component in ICL. Leveraging the fact that a two-layer transformer can implement the induction head mechanism with associative memories, we theoretically analyze the logits when a two-layer transformer is given prompts generated by a bigram model. In the experiments, we design specific prompts to evaluate whether the outputs of a two-layer transformer align with the theoretical results.
@@ 1 Introduction
In recent years, transformer-based models, such as BERT [13] and GPT [42, 43, 5], have achieved remarkable success in natural language processing. Especially, in-context learning (ICL) [14] has emerged as a groundbreaking capability within large language models (LLMs), enabling them to adapt to new tasks without traditional fine-tuning. Instead, these models leverage patterns from a prompt or input sequence, effectively learning "in context" by interpreting examples or instructions provided in real-time [33, 56]. The capability for ICL [5] in language models develops gradually during a "phase change" observed in the early stages of training. At the same time, models acquire the ability to complete complex patterns through induction heads, suggesting that induction heads are a key mechanistic basis of ICL [40].
In the context of ICL, Jiang et al. [25] discovered a phenomenon called context hijacking. This phenomenon occurs when altering the context disrupts fact recall, causing the model to generate incorrect outputs influenced by the in-context knowledge provided in the prompt. Similarly, ICL can exhibit oversight of in-context knowledge provided in the prompt, where the model relies solely on the knowledge acquired during pretraining (global knowledge) while ignoring the in-context knowledge. These types of knowledge hijacking are thought to arise from either discarding part of the prompt’s information or overly trusting the information in the prompt. Conversely, if the model can (i) comprehensively use the information in the prompt and (ii) balance in-context knowledge and global knowledge, hijacking can be prevented. In relation to the induction head, we demonstrate that employing relative positional encoding (RPE) allows the transformer to comprehensively use the information in the prompt. Furthermore, we show that even a three-layer transformer without positional encoding can achieve comprehensive use of information. Understanding how to avoid knowledge hijacking is essential for the safe and reliable use of ICL capabilities in LLMs and is a matter of significant societal importance.
In this paper, we theoretically and experimentally carry out evaluation of oversight of in-context information in a transformer, and the model’s prioritization between in-context knowledge, derived from information within the context or the prompt, and global knowledge from the training data generated from a bigram model with triggered transitions. To summarize, we make the following contributions.
  • We theoretically show that the induction head learned by a transformer using RPE can avoid oversight of in-context knowledge within the framework of associative memory.
@@ 2 Related work
Associative memory refers to the storage and retrieval of patterns, and its computational model was designed to retrieve complete memories from sufficient fragments of information in neuroscience [20, 21]. After the emergence of modern Hopfield networks [30], which store patterns in an energy function and retrieve them using an updating rule. Moreover, many studies employ the concept of associative memory in various ways. For instance, Zhang et al. [58] propose a memory unit that estimates a conditional probability distribution with Gaussian kernel smoothing. Some work considers attention blocks as memory because they read and write information flowing through the residual stream [39, 17]. Meng et al. [35] locate factual association patterns in the weight matrix of an MLP. Recently, it has become increasingly popular to regard a matrix as associative memory from a theoretical perspective, as shown in Tab. 1. Previous studies examine scaling laws in relation to model capacity [6], learning dynamics of linear model in terms of particle system [7], and the effect of low-rank approximation of weight matrix to classification involving a noisy token [8] for a linear model. Additionally, Jiang et al. [25] demonstrate that a one-layer transformer can recover latent concept from long enough context. Most relevant to our research is the work by Bietti et al. [4], which mainly discusses the learning dynamics of induction heads for a two-layer transformer, but our primary focus is on the reasoning process.
@@
Since Brown [5] introduced ICL, theoretical understanding of this phenomenon has been a major interest within the community. For example, Akyürek et al. [2] showed that transformer can implement a learning algorithm that solves linear regression. Von Oswald et al. [51] demonstrated that linear self-attention layer can emulate gradient descent. Similarly, many studies [1, 34, 59, 11] examined the theoretical connections between ICL and gradient descent. Mechanistic interpretability, which aims to understand the model behavior by reverse-engineering its internal components [37, 53, 10], is another perspective for analyzing ICL [40, 16]. In particular, Elhage et al. [16] discovered that two-layer attention-only transformers understand the previous use of the token and look for similar situations in the context. In addition, Olsson et al. [40] provided empirical evidence that induction heads are the mechanism behind ICL. Moreover, Bansal et al. [3] identified that many induction heads are also among the attention heads deemed critical for ICL. The most relevant study to ours is by Bietti et al. [4], who investigated the learning dynamics of the induction head mechanism in two-layer transformers from the perspective of associative memory. In contrast, our study focuses on the length generalization of the induction head mechanism and the influence of knowledge acquired during training and that obtained in context on the final output.
@@ 3.1 Transformer architecture
A decoder-only transformer consists of a stack of identical layers, each composed of several sub-components that enable it to generate sequential outputs in an autoregressive manner. The decoder generates sequences by predicting one token at a time, conditioned on the previously generated tokens. This process continues until an end-of-sequence token is generated or the maximum sequence length is reached. Large language models commonly adopt this kind of architecture and are trained with next token prediction. In this work, we focus on a two-layer transformer architecture analyzed by Bietti et al. [4], while we use simple relative positional encoding instead of absolute positional encoding.
@@
$x^{(1)}_{t}$ $:=\Phi_{1}x^{(0,v)}_{1:t}\sigma((W_{K}^{1}x^{(0,k)}_{1:t})^{\top}W_{Q}^{1}x^{(0,q)}_{t})$
$\qquad+x^{(0,v)}_{t}\in\mathbb{R}^{d},$ (1)
$x^{(2)}_{t}$ $:=W_{O}^{2}W_{V}^{2}x^{(1)}_{1:t}\sigma((W_{K}^{2}x_{1:t}^{(1)})^{\top}W_{Q}^{2}x_{t}^{(1)})$
$\qquad+x^{(1)}_{t}\in\mathbb{R}^{d},$ (2)
$x_{t}$ $:=W_{F}x^{(2)}_{t}+x^{(2)}_{t},$ (3)
@@
where $W_{E}\in\mathbb{R}^{d\times|\mathcal{V}|}$ is an embedding matrix, $\sigma:\mathbb{R}^{t}\to\mathbb{R}^{t}$ is the softmax function, $x^{(0,k)},x^{(0,q)},\text{ and }x^{(0,v)}$ are the token embeddings, which may or may not include positional information, $\Phi_{1}\in\mathbb{R}^{d\times d}$ represents $W_{O}^{1}W_{V}^{1}$, $W_{F}\in\mathbb{R}^{d\times d}$ is a linear projection, $W_{U}=(w_{U}(v_{1}),\dots,w_{U}(v_{|\mathcal{V}|}))^{\top}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is an unembedding matrix, and $v_{i}$ represents the $i$-th vocabulary in $\mathcal{V}$.
@@
3.2 Bigram model with triggered transitions
This section describes the bigram model with triggered tokens used in Bietti et al. [4]. In this bigram language model, transitions between certain "trigger" tokens (denoted as $q_{k}$) are modified in each sequence. Specifically, these trigger tokens are always followed by specific output tokens $o_{k}$. The model operates over a vocabulary of size $V$, with the following distributions: $\pi_{b}(\cdot\mid i)$ for bigram conditionals, $\pi_{u}$ for the unigram distribution, and $\pi_{o}(\cdot\mid i)$ to generate output tokens, along with a distribution $\pi_{q}$ for trigger tokens.
The process involves randomly sampling trigger tokens from $\pi_{q}$. Then, output tokens $o_{k}$ are sampled from $\pi_{o}$, and the sequence $z_{1:T}$ is generated as follows:
  • Sample $q_{1},\dots,q_{K}\sim\pi_{q}$, without replacement (random triggers).
  • Sample $o_{k}\sim\pi_{o}(\cdot\mid q_{k})$, with replacement.
  • Generate $z_{1}\sim\pi_{u}$ and for each subsequent token $z_{t}$ (for $t=2,\dots,T$), sample $z_{t}\mid z_{t-1}$ as follows:
    $p_{n}(j\mid i)=\begin{cases}\pi_{b}(j\mid i),&\text{if }i\notin\{q_{k}\},\\ 1\{j=o_{k}\},&\text{if }i=q_{k}.\end{cases}$
Motivation for the model
The induction head is a computational mechanism capable of detecting repeated patterns in context and facilitating the output of the tokens that complete these patterns when a trigger token appears. In written texts centered on a particular theme, it is common for identical patterns of tokens to recur multiple times. For instance, in texts about Harry Potter or machine learning, the token "Avada" is often followed by "Kedavra," and "neural" is typically followed by "network."
In a pure bigram model, the token following a given token $z$ is determined by conditional probabilities. As a result, consistent repeated patterns may not emerge, or multiple distinct patterns beginning with $z$ may appear, making it difficult to thoroughly evaluate the function of the induction head. However, by associating specific output tokens with designated trigger tokens, it becomes possible to generate multiple instances of the same pattern within the input sequence. Thus, this approach enables a more effective analysis of the induction head mechanism in scenarios that closely resemble real-world text.
@@ 3.4 Induction head
In Fig. 1, we summarize the process by which the induction head identifies and outputs repetitive patterns in the input. When the prompt $ABC\dots A$ is given, the first-layer attention uses $W_{K}^{1}$ to copy information from the previous token to the current token. Consequently, the token $B$ retains the information of $A$ in the form of $\Phi_{1}A$. Then, in the second-layer attention, the associative memory $W_{K}^{2}$ matches the pair $\Phi_{1}A$ with $A$, attending to the position of token $B$. As a result, the information of $B$ is stored in the last token $A$ as $W_{V}^{2}B$. Finally, the output matrix $W_{O}^{2}$ transforms $W_{V}^{2}B$ into $W_{U}(B)$, producing an output that follows the repetitive pattern.
@@ 4 Theoretical result
We analyze two key aspects of induction heads: the impact of positional encoding on the oversight, and the contributions of in-context information and pretrained knowledge during inference.
4.1 Oversight of in-context knowledge
In this section, we address the theoretical aspects of oversight mechanism in the induction head mechanism. In short, for transformers with APE, it can be noted that learning of the previous token head fails, resulting in prediction errors. This issue arises not only with input sequences of lengths that were encountered during training, but also with longer sequences. Formally, we observe the representation of the weight matrix $\tilde{W}_{K}^{1}$ when weight matrices are sequentially trained, in the order of $W_{O}^{2}$, $W_{K}^{2}$, and $W_{K}^{1}$, through one step of gradient descent. We refer to Theorem 3 from Bietti et al. [4]:
@@ 4.1 Oversight of in-context knowledge
$\approx\frac{\eta\alpha\alpha^{\prime}(T-1)}{T^{2}t}\left\{P(t_{q}=t-1)\left(1-\frac{1}{t}\right)\right.$
$\hskip 90.00014pt\left.+O\left(\frac{1}{V}\right)\right\},$ (6)
where $p_{t}$ is the absolute positional encoding for token position $t$.
From Eq. 1, we can see that the coefficient of $p_{t-1}p_{t}^{\top}$ which constitutes the associative memory $\tilde{W}_{K}^{1}$ is inversely proportional to $t$. Therefore, as $t$ increases – meaning as we move toward the latter part of the sequence– the function of the associative memory diminishes.
In contrast, we show that transformers employing RPE consistently direct the previous-token head to attend to the preceding token, regardless of the input sequence length. We first calculate how the transformer’s weight matrices are represented by following the same learning procedure as when using a transformer with APE. Here, we present the training outcome of the matrix $W_{K}^{1}$ as associative memory, which plays the most significant role in length generalization by constructing the previous token head. We provide the complete statement and proof in Appendix B.
@@
Under the setup described in Sec. B.2, a two-layer attention-only transformer with relative positional encoding learns the associations by one step of gradient descent for any vocabulary $v\in\mathcal{V}$:
@@
We also present the following Proposition that shows a three-layer transformer without positional encoding can implement the induction head mechanism.
@@
Here, we discuss whether to prioritize statistical information obtained from the training data or the information provided in-context. By calculating the logits that determine the output probabilities of the vocabulary in an associative memory transformer, we clarify the mechanisms through which outputs are generated.
@@
Define $U_{v}=\{u\in\mathcal{V}\mid\pi_{b}(u\mid v)>0\}$ as the support of $\pi_{b}(u\mid v)$ and $Q=\operatorname{supp}(\pi_{q})$. The induction head can be achieved by constructing matrices in the following manner.
$W_{Q}^{1}$ $=W_{Q}^{2}=I,$
$W_{K}^{1}$ $=\sum_{t=2}^{T}p_{t}p_{t-1}^{\top},$
@@
where $I$ is the identity matrix. The other matrices $W_{O}^{1}$, $W_{V}^{1}$, and $W_{O}^{2}$ are a Gaussian random matrix.
@@
$W_{K}^{1}$ $=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top},$
$W_{F}$ $=\sum_{v=1}^{V}\sum_{u=1}^{V}\log{\pi_{b}(u\mid v)}w_{U}(u)w_{E}(v)^{\top},$
Note that when the probability $\pi_{b}(u\mid v)$ approaches $+0$, its logarithm $\log{\pi_{b}(u\mid v)}$ diverges to $\infty$- ∞. If we admit $\log{\pi_{b}(u\mid v)}=-\infty$ for $u\notin U_{v}$ in the construction of $W_{F}$, the transformer does not output $u$ after the token $v$ regardless of the context. This does not align with the actual behavior of the induction head. Also, due to the limitations of floating-point representation in computers, it is not possible to represent $\infty$ directly. Thus, we consider $\epsilon$ as threshold and set $\pi_{b}(u\mid v)=\epsilon$.
@@
First, we present the following proposition, which focuses on two token patterns in the input sequence.
@@
Suppose a two-layer transformer is an associative memory transformer. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let the $t_{1}$-th token be $z_{t_{1}}=v_{1}$ following $z_{t_{1}-1}=q$, and let the $t_{2}$-th token be $z_{t_{2}}=v_{2}$ following $z_{t_{2}-1}=q$, where $v_{1}\neq v_{2}$ and $(3\leq)t_{1}<t_{2}$ without loss of generality. Also assume that there exists only one occurrence for each of $v_{1}$ and $v_{2}$ in the context $z_{1:T}$. Then, the logits $\xi_{v_{1}}$ and $\xi_{v_{2}}$ can be expressed as follows:
$\xi_{v_{1}}$ $=\exp{\left(\frac{e}{t_{1}+e-1}\right)}/Z+\log{\pi_{b}(v_{1}\mid q)},$ (7)
$\xi_{v_{2}}$ $=\exp{\left(\frac{1+e}{t_{2}+e-1}\right)}/Z+\log{\pi_{b}(v_{2}\mid q)},$ (8)
where $Z$ is the normalization factor of the softmax function.
We can observe a general trend from Eqs. 7 and 8. First, if the model contains global knowledge of both tokens $v_{1}$ and $v_{2}$, the output token is influenced by the knowledge, in addition to the in-context knowledge. This implies that the occurrence of tokens that rarely appear in the training dataset is suppressed. In our associative memory transformer model, global knowledge never promotes the occurrence of tokens but only suppresses inappropriate occurrence, because we always have $\log{\pi_{b}(\cdot\mid q)}\leq 0$. Second, if token patterns, such as $q\,v_{1}$ and $q\,v_{2}$, do not appear in the training data, global knowledge affects the token logits to the same extent, due to the way $W_{F}$ is defined using the threshold $\epsilon$ in Def. 2. This indicates that the model intensively suppresses token patterns that were not seen during training.
It is worth noting that an associative memory transformer exhibits a significant difference in scale when handling global knowledge and in-context knowledge. For instance, if $\pi_{b}(v\mid q)=0.1$, then $\log{\pi_{b}(v\mid q)}\approx-2.3$. Even with the functionality of an induction head, it is unlikely to predict the token patterns from the context. This is avoidable by reformulating $W_{O}^{2}$ as in Eq. 12 and adjusting the strength of induction head. We consider this setting in the next paragraph, along with more general input sequences.
@@
$\xi^{\prime}_{v_{2}}-\xi^{\prime}_{v_{1}}=$ $\frac{e(t_{1}-t_{2})+t_{1}+e-1}{(t_{1}+e-1)(t_{2}+e-1)}$ (9)
@@
A stronger associative memory transformer is equipped with the following weight matrices:
$W_{K}^{1}$ $=\tau_{1}\sum_{k\in Q}w_{E}(k)r_{-1}^{\top},$ (10)
$W_{K}^{2}$ $=\tau_{2}\sum_{k\in Q}w_{E}(k)(W_{O}^{1}W_{V}^{1}w_{E}(k))^{\top},$ (11)
$W_{O}^{2}$ $=\tau_{3}\sum_{v=1}^{V}w_{U}(v)(W_{V}^{2}w_{E}(v))^{\top},$ (12)
@@
Suppose a two-layer transformer is a stronger associative memory transformer as defined in Eqs. 10, 11, and 12. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let $f(v)$ be the number of token pattern "$q\,v$" appearing in the sequence for vocabulary $v\in\mathcal{V}$. For sufficiently large $\tau_{1}$ and $\tau_{2}$, the logits $\xi_{v}$ can be expressed as follows:

### v2

# Rethinking Associative Memory Mechanism in Induction Head

@@ Abstract
Induction head mechanism is a part of the computational circuits for in-context learning (ICL) that enable large language models (LLMs) to adapt to new tasks without fine-tuning. Most existing work explains the training dynamics behind acquiring such a powerful mechanism. However, the model’s ability to coordinate in-context information over long contexts and global knowledge acquired during pretraining remains poorly understood. This paper investigates how a two-layer transformer thoroughly captures in-context information and balances it with pretrained bigram knowledge in next token prediction, from the viewpoint of associative memory. We theoretically analyze the representation of weight matrices in attention layers and the resulting logits when a transformer is given prompts generated by a bigram model. In the experiments, we design specific prompts to evaluate whether the outputs of the trained transformer align with the theoretical results.
@@ 1 Introduction
In recent years, transformer-based models, such as BERT [14] and GPT [44, 45, 6], have achieved remarkable success in natural language processing. Especially, in-context learning (ICL) [6, 15] has emerged as a groundbreaking capability within large language models (LLMs), enabling them to adapt to new tasks without traditional fine-tuning. Instead, these models leverage patterns from a prompt or input sequence, effectively learning "in context" by interpreting examples or instructions provided in real-time [35, 58]. During a phase change, models acquire the ability to complete complex patterns through induction heads, suggesting that induction heads are a key mechanistic basis of ICL [42]. It is known that the induction head mechanism can emerge in two-layer transformers [17], and many theoretical studies have focused on analyzing two-layer architectures to understand how induction head develops [16, 41, 10]. Among them, Bietti et al. [5] demonstrated, both theoretically and empirically, that training a transformer leads to weight matrices behaving as associative memories, composing the induction head mechanism.
However, the effectiveness of the induction head mechanism tends to diminish for the latter part of a sequence [5], which does not align with the behavior observed in real-world LLM. In addition, while the learning dynamics of the induction head have been elucidated in the existing work, it remains unclear how in-context knowledge provided in the prompt and global knowledge acquired during pretraining influence the output of two-layer transformers. Understanding how a transformer predicts the next token is crucial, as it allows us to model phenomena such as context hijacking [27], in which altering the context disrupts factual recall, causing the model to produce incorrect outputs influenced by misleading in-context information.
In this paper, we theoretically and empirically show, from the perspective of associative memory, how a two-layer transformer with relative positional encoding (RPE) can avoid failure in detecting patterns that appear in the latter part of a sequence. We also investigate how the model prioritizes in-context knowledge versus global knowledge when predicting the next token, using training data generated by a bigram model with triggered transitions. To summarize, we make the following contributions:
  • We theoretically show that the induction head learned by a transformer using RPE can avoid oversight of patterns within the framework of associative memory.
@@ 2 Related work
Associative memory refers to the storage and retrieval of patterns, and its computational model was designed to retrieve complete memories from sufficient fragments of information in neuroscience [3, 22, 23]. After the emergence of modern Hopfield networks [32], which store patterns in an energy function and retrieve them using an updating rule. Moreover, many studies employ the concept of associative memory in various ways. For instance, Zhang et al. [60] propose a memory unit that estimates a conditional probability distribution with Gaussian kernel smoothing. Some work considers attention blocks as memory because they read and write information flowing through the residual stream [41, 18]. Meng et al. [37] locate factual association patterns in the weight matrix of an MLP. Recently, it has become increasingly popular to regard a matrix as associative memory from a theoretical perspective, as shown in Tab. 1. Previous studies examine scaling laws in relation to model capacity [7], learning dynamics of linear model in terms of particle system [8], and the effect of low-rank approximation of weight matrix to classification involving a noisy token [9] for a linear model. Additionally, Jiang et al. [27] demonstrate that a one-layer transformer can recover latent concept from long enough context.
@@
Since Brown [6] introduced ICL, theoretical understanding of this phenomenon has been a major interest within the community. For example, Akyürek et al. [2] showed that transformer can implement a learning algorithm that solves linear regression. Von Oswald et al. [53] demonstrated that linear self-attention layer can emulate gradient descent. Similarly, many studies [1, 36, 61, 12] examined the theoretical connections between ICL and gradient descent. Mechanistic interpretability, which aims to understand the model behavior by reverse-engineering its internal components [39, 55, 11], is another perspective for analyzing ICL [42, 17]. In particular, Elhage et al. [17] discovered that two-layer attention-only transformers understand the previous use of the token and look for similar situations in the context. In addition, Olsson et al. [42] provided empirical evidence that induction heads are the mechanism behind ICL. Moreover, Bansal et al. [4] identified that many induction heads are also among the attention heads deemed critical for ICL. The most relevant study to ours is by Bietti et al. [5], who investigated the learning dynamics of the induction head mechanism in two-layer transformers from the perspective of associative memory. In contrast, our study focuses on the functionality of the induction head mechanism and the influence of knowledge acquired during training and that obtained in context on the final output.
@@ 3.1 Transformer architecture
In this work, we focus on a two-layer transformer described below.
@@
$x^{(1)}_{t}$ $:=\Phi_{1}x^{(0,v)}_{1:t}\sigma((W_{K}^{1}x^{(0,k)}_{1:t})^{\top}W_{Q}^{1}x^{(0,q)}_{t})+x^{(0,v)}_{t}$ (1)
$\in\mathbb{R}^{d},$
$x^{(2)}_{t}$ $:=W_{O}^{2}W_{V}^{2}x^{(1)}_{1:t}\sigma((W_{K}^{2}x_{1:t}^{(1)})^{\top}W_{Q}^{2}x_{t}^{(1)})+x^{(1)}_{t}$ (2)
$\in\mathbb{R}^{d},$
$x_{t}$ $:=W_{2}(\operatorname{ReLU}(W_{1}x^{(2)}_{t}))+x^{(2)}_{t},$ (3)
@@
where $W_{E}\in\mathbb{R}^{d\times|\mathcal{V}|}$ is an embedding matrix, $\sigma:\mathbb{R}^{t}\to\mathbb{R}^{t}$ is the softmax function, $x^{(0,k)},x^{(0,q)},\text{ and }x^{(0,v)}$ are the token embeddings, which may or may not include positional information, $\Phi_{1}\in\mathbb{R}^{d\times d}$ represents $W_{O}^{1}W_{V}^{1}$, $W_{U}=(w_{U}(v_{1}),\dots,w_{U}(v_{|\mathcal{V}|}))^{\top}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is an unembedding matrix, and $v_{i}$ represents the $i$-th vocabulary in $\mathcal{V}$.
The feed-forward layer consists of matrices $W_{1}\in\mathbb{R}^{d\times V}$ and $W_{2}\in\mathbb{R}^{V\times d}$.
@@ 3.3 Induction head
In Fig. 2, we summarize the process by which the induction head identifies and outputs repetitive patterns in the input. When the prompt $ABC\dots A$ is given, the first-layer attention, which is called previous token head, copies information from the previous token to the current token. Consequently, the token $B$ retains the information of $A$ in the form of $\Phi_{1}A$. Then, in the second-layer attention, the associative memory $W_{K}^{2}$ matches the pair $\Phi_{1}A$ with $A$, attending to the position of token $B$. As a result, the information of $B$ is stored in the last token $A$ as $W_{V}^{2}B$. Finally, the output matrix $W_{O}^{2}$ transforms $W_{V}^{2}B$ into $W_{U}(B)$, producing an output that follows the repetitive pattern.
3.4 Bigram model with triggered transitions
This section describes the bigram model with triggered tokens used in Bietti et al. [5].
In this bigram language model over a vocabulary of size $V$, we first prepare trigger tokens $q_{k}$ sampled from a distribution $\pi_{q}$ and output tokens $o_{k}$ sampled from $\pi_{o}$ that always follow their corresponding trigger token in every bigram sequence. Then, the sequence $z_{1:T}$ is generated accordingly to the bigram conditionals $\pi_{b}(\cdot\mid\cdot)$, but when a trigger token is generated, the next token is guaranteed to be its corresponding output token:
$\begin{cases}\pi_{b}(j\mid i)&\text{if }i\notin\{q_{k}\},\\ 1\{j=o_{k}\}&\text{if }i=q_{k}.\end{cases}$
Motivation for the model
In written texts centered on a particular theme, it is common for identical patterns of tokens to recur multiple times. For instance, in texts about Harry Potter or machine learning, the token "Avada" is often followed by "Kedavra," and "neural" is typically followed by "network." The induction head is a computational mechanism capable of detecting repeated patterns in context and facilitating the output of the tokens that complete these patterns when a trigger token appears.
The next token following a given token $z$ is determined by conditional probabilities in a pure bigram model. As a result, bigram sequences often lack consistent repeated patterns, which prevents the induction head from emerging. In contrast, pairs of trigger and output tokens in a sequence enable two-layer transformers to learn a circuit that leverages repeated patterns, rather than relying solely on bigram conditionals.
@@ 4 Theoretical result
We analyze two key aspects of induction heads: the impact of positional encoding on the oversight of patterns in a sequence, and the contributions of in-context information and pretrained knowledge during inference.
4.1 Oversight of patterns
In this section, we address the theoretical aspects of overlooking patterns in a sequence by the induction head mechanism.
We argue that the learning of the previous token head fails for transformers with APE, and thus, the model cannot take patterns in the latter half of a sequence into account for the prediction. Mathematically, we observe the representation of the weight matrix $\tilde{W}_{K}^{1}$ when weight matrices are sequentially trained, in the order of $W_{O}^{2}$, $W_{K}^{2}$, and $W_{K}^{1}$, through one step of gradient descent. We refer to Theorem 3 from Bietti et al. [5]:
@@ 4.1 Oversight of patterns
$\approx\frac{\eta\alpha\alpha^{\prime}(T-1)}{T^{2}t}$
×$\quad\times\left\{\mathbb{P}(t_{q}=t-1)\left(1-\frac{1}{t}\right)\right.\left.+O\left(\frac{1}{V}\right)\right\},$ (6)
where $p_{t}$ is the absolute positional encoding for token position $t$, $\alpha$, $\alpha^{\prime}\in\mathbb{R}$ are constants, and $\mathbb{P}(t_{q}=t-1)$ denote the probability that the first trigger token appears at position $t-1$.
From Eq. 1, we can see that the score for the pair of $p_{t-1}$ and $p_{t}$ is inversely proportional to $t$. Therefore, as $t$ increases – meaning as we move toward the latter part of the sequence – the function of the associative memory diminishes.
In contrast, we prove that transformers employing RPE consistently direct the previous token head to attend to the preceding token, regardless of the input sequence length. We calculate how the transformer’s weight matrices are represented by following the same learning procedure as when using a transformer with APE. We again present the training outcome of the matrix $W_{K}^{1}$ as associative memory, which is the key component for the previous token head. We provide the complete statement and proof in Appendix C.
@@
Under the setup described in Sec. C.2, a two-layer attention-only transformer with relative positional encoding learns the associations by one step of gradient descent for any vocabulary $v\in\mathcal{V}$:
@@
We have seen that positional encoding influences the behavior of induction head, but we also show that a three-layer transformer without positional encoding can implement the induction head mechanism.
@@
Here, we discuss how a two-layer transformer uses statistical information obtained from the training data and the information provided in-context by calculating the logits.
@@
Define $Q$ as the support of $\pi_{q}$. The induction head can be achieved by constructing matrices in the following manner.
$W_{Q}^{1}=W_{Q}^{2}=I,\qquad W_{K}^{1}=\sum_{t=2}^{T}p_{t}p_{t-1}^{\top},$
@@
where $I$ is the identity matrix, and the matrices $W_{O}^{1}$, $W_{V}^{1}$, and $W_{O}^{2}$ are a Gaussian random matrix.
@@
$W_{K}^{1}=\sum_{k\in Q}w_{E}(k)r_{-1}^{\top},\qquad W_{1}=\begin{pmatrix}w_{E}(v_{1})^{\top}\\ w_{E}(v_{2})^{\top}\\ \vdots\\ w_{E}(v_{V})^{\top}\end{pmatrix},$
$W_{2}=\begin{pmatrix}\sum_{u=1}^{V}\log{\pi_{b}(u\mid v_{1})}w_{U}(u)^{\top}\\ \vdots\\ \sum_{u=1}^{V}\log{\pi_{b}(u\mid v_{V})}w_{U}(u)^{\top}\end{pmatrix}^{\top}.$
This construction reveals that the feed-forward layer in Eq. 3 of Sec. 3.1 functions explicitly as a key-value memory [19]. Each row of $W_{1}$ (key) acts as a detector for an input pattern $w_{E}(v)$, and each column of $W_{2}$ (value) encodes global knowledge, i.e., the distribution $\pi_{b}$ over the output vocabulary. Appendix B discusses the difference from the architecture of Bietti et al. [5]. Note that when the probability $\pi_{b}(u\mid v)$ approaches $+0$, its logarithm $\log{\pi_{b}(u\mid v)}$ diverges to $\infty$- ∞. If we admit $\log{\pi_{b}(u\mid v)}=-\infty$, the transformer never outputs $u$ regardless of context. This does not align with the actual behavior of the induction head. Also, due to the limitations of floating-point representation in computers, it is not possible to represent $\infty$ directly. Thus, we consider $\epsilon$ as threshold and set $\pi_{b}(u\mid v)=\epsilon$.
@@
We present the following proposition, which focuses on two token patterns in the input sequence.
@@
Suppose a two-layer transformer is an associative memory transformer. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let the $t_{1}$-th token be $z_{t_{1}}=v_{1}$ following $z_{t_{1}-1}=q$, and let the $t_{2}$-th token be $z_{t_{2}}=v_{2}$ following $z_{t_{2}-1}=q$, where $v_{1}\neq v_{2}$ and $(3\leq)t_{1}<t_{2}$ without loss of generality.
Also assume that there exists only one occurrence for each of $v_{1}$ and $v_{2}$ in the context $z_{1:T}$. Then, the logits $\xi_{v_{1}}$ and $\xi_{v_{2}}$ can be expressed as follows:
$\xi_{v_{1}}$ $=\frac{e^{\frac{e}{t_{1}+e-1}}}{Z}+\log{\pi_{b}(v_{1}\mid q)},$
$\xi_{v_{2}}$ $=\frac{e^{\frac{1+e}{t_{2}+e-1}}}{Z}+\log{\pi_{b}(v_{2}\mid q)},$ (7)
where $Z$ is the normalization factor of the softmax function and $e$ is Euler’s number.
We can observe a general trend from Eqs. 2. First, if the model contains global knowledge of both tokens $v_{1}$ and $v_{2}$, the output token is influenced by the knowledge, in addition to the in-context knowledge. In particular, the logits of tokens that rarely appear in the training dataset become significantly low. In our associative memory transformer model, global knowledge never promotes the output of any specific token; rather, it solely suppresses inappropriate predictions, because we always have $\log{\pi_{b}(\cdot\mid q)}\leq 0$. Second, if token patterns, such as $q\,v_{1}$ and $q\,v_{2}$, do not appear in the training data, global knowledge affects the token logits to the same extent, due to the way $W_{F}$ is defined using the threshold $\epsilon$ in Def. 2. It is worth noting that an associative memory transformer exhibits a significant difference in scale when handling global knowledge and in-context knowledge. For instance, if $\pi_{b}(v\mid q)=0.1$, then $\log{\pi_{b}(v\mid q)}\approx-2.3$. Even with the functionality of an induction head, it is unlikely to predict the token patterns from the context. This is avoidable by reformulating $W_{O}^{2}$ as in Def. 3 and adjusting the strength of induction head. We consider this setting in the next paragraph, along with more general input sequences.
@@
$\xi^{\prime}_{v_{2}}-\xi^{\prime}_{v_{1}}=$ $\frac{e(t_{1}-t_{2})+t_{1}+e-1}{(t_{1}+e-1)(t_{2}+e-1)}.$ (8)
@@
We redefine the following matrices $W_{K}^{1},W_{K}^{2}$ and $W_{O}^{2}$ in Def. 2 for a stronger associative memory transformer
$\hat{W}_{K}^{1}=\tau_{1}W_{K}^{1},\quad\hat{W}_{K}^{2}=\tau_{2}W_{K}^{2},\quad\hat{W}_{O}^{2}=\tau_{3}W_{O}^{2},$ (9)
@@
Suppose a two-layer transformer is a stronger associative memory transformer. Given a length-$T$ sequence $z_{1:T}$ where the last token is $z_{T}=q$, let $f(v)$ be the number of token pattern "$q\,v$" appearing in the sequence for vocabulary $v\in\mathcal{V}$. For sufficiently large $\tau_{1}$ and $\tau_{2}$, the logits $\xi_{v}$ can be expressed as follows: